# Complete non-ambiguous trees and associated permutations: new enumerative results

### v2

@@ 4.3. Permutations with $5$ CNATs
Let $G$ be a graph. If $t_{G}=5$, then the $2$-core $\mathrm{Core}_{2}\left([\right)G]$ is two-connected.
@@
Let $G^{\prime}:=\mathrm{Core}_{2}\left([\right)G]$. Suppose that $G^{\prime}$ has a cut-vertex $v$ (i.e. removing $v$ disconnects the graph), and let $G^{\prime}_{1},G^{\prime}_{2},\cdots,G^{\prime}_{k}$ be the components of $G^{\prime}$ joined at $v$ (for some $k\geq 2$). By [24, Property (iv)], we have $t_{G^{\prime}}=\prod\limits_{1\leq i\leq k}t_{G^{\prime}_{i}}$. But since $t_{G^{\prime}}=t_{G}=5$ is prime, this implies that there must exist $i$ such that $t_{G_{i}}=1$ (in fact all but one of the components must satisfy this). Lemma 2.4 then implies that $G^{\prime}_{i}$ must be a tree, which contradicts the fact that $G^{\prime}$ is the $2$-core of a graph. ∎
@@
Let $C$ be a $4$-cycle in $G$. Consider a vertex $v\notin V(C)$. Since $G$ is two-connected, there must exist two paths from $v$ to distinct vertices of $C$ which don’t intersect (except at $v$). Let us denote $P_{1}$ and $P_{2}$ these two paths, ending respectively at vertices $v_{1}$ and $v_{2}$ of $C$ (with $v_{1}\neq v_{2}$). If $v_{1}$ and $v_{2}$ are adjacent vertices in $C$, then $C\setminus\{(v_{1},v_{2})\}\cup P_{1}\cup P_{2}$ would form a cycle of length at least $5$, a contradiction. Therefore $v_{1}$ and $v_{2}$ must be diagonally opposite vertices in $C$. Taking the union of a half-cycle from $v_{1}$ to $v_{2}$ together with $P_{1}$ and $P_{2}$ then gives a cycle of length $2+|E(P_{1})|+|E(P_{2})|$. Since all cycles in $G$ have length at most $4$, this implies that $P_{1}$ and $P_{2}$ each consist of a single edge, and the union of these with $C$ gives the desired subgraph $H$. ∎
@@
Let $G$ be a simple graph. Then $t_{G}=5$ if, and only if, $G$ is a decorated $6$-cycle.
@@
If $G$ is a decorated $6$-cycle, then we have $t_{G}=t_{\mathrm{Core}_{2}\left([\right)G]}=6-1=5$ by Lemma 2.6. Suppose now that $G$ is a simple graph such that $t_{G}=5$, and define $G^{\prime}:=\mathrm{Core}_{2}\left([\right)G]$ to be the $2$-core of $G$. We wish to show that $G^{\prime}$ is (isomorphic to) the $6$-cycle $C_{6}$. Let $k$ be the length of the longest cycle in $G^{\prime}$ (such a cycle must exist since otherwise we would have $t_{G}=t_{G^{\prime}}=1$ by Lemma 2.4). That is, $k$ is the maximal value such that $G^{\prime}$ contains a $k$-cycle $v_{0},v_{1},\cdots,v_{k-1},v_{k}=v_{0}$ where the first $k-1$ vertices are distinct. With a slight abuse of notation, we denote this $k$-cycle $C_{k}$.
We first note that we must have $k\leq 6$. Indeed, otherwise we would have $t_{G}=t_{G^{\prime}}\geq t_{C_{k}}=k-1\geq 7-1=6$, a contradiction (using Lemmas 2.6 and 2.8). Moreover, if $k=6$, then $G^{\prime}$ must be the $6$-cycle. Otherwise, $G$ would contain a $6$-cycle to which its $2$-core is not isomorphic, which implies $t_{G}>t_{C_{6}}=5$ by Lemma 2.8.
It therefore remains to show that we cannot have $k\leq 5$. First, note that $G^{\prime}$ cannot be the $k$-cycle $C_{k}$ in this case, since otherwise we would have $t_{G}=t_{G^{\prime}}=t_{C_{k}}=k-1\leq 4$, a contradiction. A chord of $C_{k}$ in $G^{\prime}$ is a path in $G^{\prime}$ between two distinct vertices $v_{i},v_{j}$ of $C_{k}$ which only intersects $C_{k}$ at the two end-points $v_{i}$ and $v_{j}$. We claim that $C_{k}$ must contain a chord in $G^{\prime}$. Indeed, since $G^{\prime}$ is not $C_{k}$, then there must be an edge $(v,w)$ for some vertex $v\in C_{k}$ such that $(v,w)$ is not an edge of $C_{k}$ (i.e. if $v=v_{i}$, then $w\notin\{v_{i-1},v_{i+1}\}$). If $w\in C_{k}$, then the edge $(v,w)$ is a chord by definition, so suppose that $w\notin C_{k}$. Since $G^{\prime}$ is two-connected by Lemma 4.19, removing $v$ does not disconnect the graph, so there must be a path from $w$ to one of the neighbours $v^{\prime}$ of $v$ on the cycle $C_{k}$ which does not use the edge $(v,v^{\prime})$. Taking this path from $w$ to the first point at which it intersects the cycle $C_{k}$ gives the desired chord. We now proceed on a case-by-case basis on the value of $k$. In each case, the acyclic orientation numbers can be checked e.g. through applying Equation (1) or through using the SageMath software [23].
Case $k=3$.:
Since $G^{\prime}$ is simple, any chord of $C_{3}$ must have length at least $2$ (otherwise it would be a multiple edge), let us say that this chord is $v_{0},w_{1},\cdots,w_{k},v_{1}$, with $k\geq 1$ and $w_{i}\notin\{v_{0},v_{1},v_{2}\}$ for $1\leq i\leq k$. Then $v_{0},w_{1},\cdots,w_{k},v_{1},v_{2},v_{0}$ is a cycle of length at least $4$ in $G^{\prime}$, a contradiction.
Case $k=4$.:
@@
    If $|V(G^{\prime})|\geq 5$, we apply Lemma 4.21. One can check that the graph $H$ of Figure 17 satisfies $t_{H}=7$, and by Lemma 2.8 we then have $t_{G^{\prime}}\geq t_{H}$, which contradicts $t_{G^{\prime}}=5$.
    If $|V(G^{\prime})|=4$, then $G^{\prime}$ must be $C_{4}$ with either one or two chords of length $1$, as in Figure 18. One can check that these have acyclic orientation numbers $4$ and $6$ respectively, neither of which equals $5$.
@@
Case $k=5$.:
[…]
@@
Suppose that $G$ is such that $t_{G}=5$. By Proposition 4.22, its $2$-core must be isomorphic to the $6$-cycle. In particular, the original graph $G$ induces a $6$-cycle (the $2$-core of $G$ is always an induced subgraph of $G$). But from Point (3) of Proposition 2.10 this is impossible if $G$ were a permutation graph. As such, we have shown that a graph $G$ satisfying $t_{G}=5$ cannot be a permutation graph, which implies that there are no permutations with $5$ CNATs by Theorem 2.12, as desired. ∎
Problem 4.23.
Our proof of Theorem 4.18 relies heavily on the structure of permutation graphs, and the fact that $\mathrm{cnat}\left([\right)\pi]=t_{G_{\pi}}$ (Theorem 2.12), which is non-trivial. It would be interesting to find a more direct combinatorial proof of this result.
@@ 4.4. Permutations with maximal numbers of CNATs
For any $n\geq 1$, we have $\max\{\mathrm{cnat}\left([\right)\pi];\pi\in S_{n}\}=(n-1)!$, and this maximum is achieved only for the decreasing permutation $\mathrm{dec}_{[}n]=n(n-1)\cdots 1$. In other words, we have $B(n,(n-1)!)=\{\mathrm{dec}_{[}n]\}$ and $B(n,k)=\emptyset$ for all $k>(n-1)!$.
@@
That $\mathrm{cnat}\left([\right){\mathrm{dec}_{[}n]}]=(n-1)!$ is a consequence of Theorem 3.5. We therefore need to show that if $\pi\neq\mathrm{dec}_{[}n]$, we have $\mathrm{cnat}\left([\right)\pi]<\mathrm{cnat}\left([\right){\mathrm{dec}_{[}n]}]$. For this, note that the permutation graph $G_{\mathrm{dec}_{[}n]}$ associated with the decreasing permutation is the complete graph $K_{n}$ on $n$ vertices, and that $\mathrm{dec}_{[}n]$ is the only permutation whose graph is the complete graph. Now for $\pi\neq\mathrm{dec}_{[}n]$, the permutation graph $G_{\pi}$ is a graph on $n$ vertices which is not complete. The complete graph $K_{n}$ can therefore be obtained from $G_{\pi}$ through a non-empty sequence of edge additions. By Lemma 2.7, this implies that $t_{G_{\pi}}<t_{K_{n}}=t_{G_{\mathrm{dec}_{[}n]}}$. The result then follows from Theorem 2.12. ∎
@@
[…]
@@ 5. Conclusion and future perspectives
In this paper, we deepened the combinatorial studies of CNATs initiated in the literature (see e.g.  [3, 11, 2]) by exploiting a connection between CNATs with a given permutation and the Tutte polynomial of the corresponding permutation graph that appeared (in equivalent form) in [13]. In Theorem 3.5, we provided a new bijection between permutations and so-called upper-diagonal CNATs (CNATs whose associated permutation is the decreasing permutation $n(n-1)\cdots 1$), which has the added benefit of preserving certain statistics of these objects. This bijection is defined recursively, based on two operations on labelled CNATs called top-row decomposition and top-row deletion.
We then investigated the enumeration of permutations with a given number of CNATs, answering a number of conjectures from [11]. We gave separate characterisations of permutations with exactly $1$, $2$ or $3$ CNATs in terms of so-called quadrants of the permutation’s plot, and in terms of permutation patterns. This allowed us to establish bijections between marked permutations with a single CNAT and permutations with $2$ CNATs (Theorem 4.3), and between permutations with $2$ CNATs and permutations with $3$ CNATs (Theorem 4.16). From this we obtained enumerative formulas for permutations with $2$ or $3$ CNATs (Corollaries 4.5 and 4.17). Finally, we showed that there are no permutations of any length with exactly $5$ CNATs (Theorem 4.18), and that the maximal number of CNATs associated with a given permutation of length $n$ is $(n-1)!$, achieved uniquely for the decreasing permutation $n(n-1)\cdots 1$.
@@
  • Find a more “direct” combinatorial proof of the fact that there are no permutations with exactly $5$ CNATs, as explained in Problem 4.23.
  • Investigate the enumerative sequences $\big{(}b(n,k)\big{)}_{n\geq 1}$ for more values of $k$. In this paper, we have looked at the cases $k=1,2,3,5,(n-1)!$. The paper [11] suggests that the entries for $k=4,6,7$ do indeed appear in the OEIS [17]. As explained in Remark 4.25, this is not the case for general factorial values of $k$, so it may be that these specific values essentially rely on there being relatively few permutation graphs with the specified acyclic orientation number. For instance, it seems plausible that the only graphs with acyclic orientation number equal to $4$ are the two graphs made up of two triangles in Figures 13 and 17(a) with some trees attached. It also would be particularly interesting to know if there are values of $k$ other than $5$ for which $b(n,k)$ is always 00.
@@
  • A final direction of future research could be to restrict ourselves to certain subsets of permutations, for example derangements (permutations with no fixed point). Fixed points seem to play an important role in the enumeration of permutations according to their number of associated CNATs, since a fixed point implies the existence of a $321$ pattern, which in turn means a $3$-cycle in the permutation graph (and cycles play an important role in graph orientations). For instance, we know that there are no derangements with $2$ CNATs (Proposition 4.10), while permutations with $1$ and $3$ CNATs are all derangements (Propositions 4.2 and 4.15). It is intriguing to consider other values of $k$ and check whether there are derangements or not which have $k$ CNATs. Besides derangements, we could also restrict ourselves to permutations containing or avoiding certain patterns.

### v4

@@ 4.3. Permutations with $5$ CNATs
Let $G$ be a graph. If $a_{G}=5$, then the $2$-core $\mathrm{Core}_{2}\left([\right)G]$ is two-connected.
@@
Let $G^{\prime}:=\mathrm{Core}_{2}\left([\right)G]$. Suppose that $G^{\prime}$ has a cut-vertex $v$ (i.e. removing $v$ disconnects the graph), and let $G^{\prime}_{1},G^{\prime}_{2},\ldots,G^{\prime}_{k}$ be the components of $G^{\prime}$ joined at $v$ (for some $k\geq 2$). By [24, Property (iv)], we have $a_{G^{\prime}}=\prod\limits_{1\leq i\leq k}a_{G^{\prime}_{i}}$. But since $a_{G^{\prime}}=a_{G}=5$ is prime, this implies that there must exist $i$ such that $a_{G_{i}}=1$ (in fact all but one of the components must satisfy this). Lemma 2.4 then implies that $G^{\prime}_{i}$ must be a tree, which contradicts the fact that $G^{\prime}$ is the $2$-core of a graph. ∎
@@
Let $C$ be a $4$-cycle in $G$. Consider a vertex $v\notin V(C)$. Since $G$ is two-connected, there must exist two paths from $v$ to distinct vertices of $C$ which do not intersect (except at $v$). Let us denote $P_{1}$ and $P_{2}$ these two paths, ending respectively at vertices $v_{1}$ and $v_{2}$ of $C$ (with $v_{1}\neq v_{2}$). If $v_{1}$ and $v_{2}$ are adjacent vertices in $C$, then $C\setminus\{(v_{1},v_{2})\}\cup P_{1}\cup P_{2}$ would form a cycle of length at least $5$, a contradiction. Therefore $v_{1}$ and $v_{2}$ must be diagonally opposite vertices in $C$. Taking the union of a half-cycle from $v_{1}$ to $v_{2}$ together with $P_{1}$ and $P_{2}$ then gives a cycle of length $2+|E(P_{1})|+|E(P_{2})|$. Since all cycles in $G$ have length at most $4$, this implies that $P_{1}$ and $P_{2}$ each consist of a single edge, and the union of these with $C$ gives the desired subgraph $H$. ∎
@@
Let $G$ be a simple graph. Then $a_{G}=5$ if, and only if, $G$ is a decorated $6$-cycle.
@@
If $G$ is a decorated $6$-cycle, then we have $a_{G}=a_{\mathrm{Core}_{2}\left([\right)G]}=6-1=5$ by Lemma 2.6. Suppose now that $G$ is a simple graph such that $a_{G}=5$, and define $G^{\prime}:=\mathrm{Core}_{2}\left([\right)G]$ to be the $2$-core of $G$. We wish to show that $G^{\prime}$ is (isomorphic to) the $6$-cycle $C_{6}$. Let $k$ be the length of the longest cycle in $G^{\prime}$ (such a cycle must exist since otherwise we would have $a_{G}=a_{G^{\prime}}=1$ by Lemma 2.4). That is, $k$ is the maximal value such that $G^{\prime}$ contains a $k$-cycle $v_{0},v_{1},\ldots,v_{k-1},v_{k}=v_{0}$ where the first k1k-1 vertices are distinct. With a slight abuse of notation, we denote this $k$-cycle $C_{k}$.
We first note that we must have $k\leq 6$. Indeed, otherwise we would have $a_{G}=a_{G^{\prime}}\geq a_{C_{k}}=k-1\geq 7-1=6$, a contradiction (using Lemmas 2.6 and 2.8). Moreover, if $k=6$, then $G^{\prime}$ must be the $6$-cycle. Otherwise, $G$ would contain a $6$-cycle to which its $2$-core is not isomorphic, which implies $a_{G}>a_{C_{6}}=5$ by Lemma 2.8.
It therefore remains to show that we cannot have $k\leq 5$. First, note that $G^{\prime}$ cannot be the $k$-cycle $C_{k}$ in this case, since otherwise we would have $a_{G}=a_{G^{\prime}}=a_{C_{k}}=k-1\leq 4$, a contradiction. A chord of $C_{k}$ in $G^{\prime}$ is a path in $G^{\prime}$ between two distinct vertices $v_{i},v_{j}$ of $C_{k}$ which only intersects $C_{k}$ at the two end-points $v_{i}$ and $v_{j}$. We claim that $C_{k}$ must contain a chord in $G^{\prime}$. Indeed, since $G^{\prime}$ is not $C_{k}$, then there must be an edge $(v,w)$ for some vertex $v\in C_{k}$ such that $(v,w)$ is not an edge of $C_{k}$ (i.e. if $v=v_{i}$, then $w\notin\{v_{i-1},v_{i+1}\}$). If $w\in C_{k}$, then the edge $(v,w)$ is a chord by definition, so suppose that $w\notin C_{k}$. Since $G^{\prime}$ is two-connected by Lemma 4.18, removing $v$ does not disconnect the graph, so there must be a path from $w$ to one of the neighbours $v^{\prime}$ of $v$ on the cycle $C_{k}$ which does not use the edge $(v,v^{\prime})$. Taking the edge $(v,w)$ together with this path from $w$ to the first point at which it intersects the cycle $C_{k}$ gives the desired chord. We now proceed on a case-by-case basis on the value of $k$. In each case, the acyclic orientation numbers can be checked e.g. through applying Equation (1) or through using the SageMath software [23].
Case $k=3$:
Since $G^{\prime}$ is simple, any chord of $C_{3}$ must have length at least $2$ (otherwise it would be a multiple edge), let us say that this chord is $v_{0},w_{1},\ldots,w_{k},v_{1}$, with $k\geq 1$ and $w_{i}\notin\{v_{0},v_{1},v_{2}\}$ for $1\leq i\leq k$. Then $v_{0},w_{1},\ldots,w_{k},v_{1},v_{2},v_{0}$ is a cycle of length at least $4$ in $G^{\prime}$, a contradiction.
Case $k=4$:
@@
    If $|V(G^{\prime})|\geq 5$, we apply Lemma 4.20. One can check that the graph $H$ of Figure 17 satisfies $t_{H}=7$, and by Lemma 2.8 we then have $a_{G^{\prime}}\geq t_{H}$, which contradicts $a_{G^{\prime}}=5$.
    If $|V(G^{\prime})|=4$, then $G^{\prime}$ must be $C_{4}$ with either one or two chords of length 11, as in Figure 18. One can check that these have acyclic orientation numbers $4$ and $6$ respectively, neither of which equals $5$.
@@
Case $k=5$:
Recall that $C_{5}$ must contain a chord. This chord should have length $1$ or $2$ (otherwise we would get a cycle of length at least $6$), meaning that $G$ contains (as a subgraph) one of the two cases of Figure 19. One can check that the graph $H_{1}$ satisfies $a_{H_{1}}=6$, while the graph $H_{2}$ yields $a_{H_{2}}=10$. [^4] Contracting a chord edge gives $H_{1}$, while deleting it gives a decorated $5$-cycle, so $a_{H_{2}}=a_{H_{1}}+a_{C_{5}}=6+4=10$. In particular, Lemma 2.8 implies that $a_{G^{\prime}}\geq\min\{a_{H_{1}},a_{H_{2}}\}=6$, contradicting $a_{G^{\prime}}=5$. This completes the proof of the proposition.
@@
Suppose that $G$ is such that $a_{G}=5$. By Proposition 4.21, its $2$-core must be isomorphic to the $6$-cycle. In particular, the original graph $G$ induces a $6$-cycle (the $2$-core of $G$ is always an induced subgraph of $G$). But from Point (3) of Proposition 2.10 this is impossible if $G$ were a permutation graph. As such, we have shown that a graph $G$ satisfying $a_{G}=5$ cannot be a permutation graph, which implies that there are no permutations with $5$ CNATs by Theorem 2.13, as desired. ∎
Problem 4.22.
Our proof of Theorem 4.17 relies heavily on the structure of permutation graphs, and the fact that $\mathrm{cnat}\left([\right)\pi]=a_{G_{\pi}}$ (Theorem 2.13), which is non-trivial. It would be interesting to find a more direct combinatorial proof of this result.
@@ 4.4. Permutations with maximal numbers of CNATs
For any $n\geq 1$, we have $\max\{\mathrm{cnat}\left([\right)\pi];\pi\in S_{n}\}=(n-1)!$, and this maximum is achieved only for the decreasing permutation $\mathrm{dec}_{[}n]=n(n-1)\cdots 1$. In other words, we have $B(n,(n-1)!)=\{\mathrm{dec}_{[}n]\}$ and $B(n,k)=\emptyset$ for all k>(n1)!k>(n-1)!.
@@
That $\mathrm{cnat}\left([\right){\mathrm{dec}_{[}n]}]=(n-1)!$ is a consequence of Theorem 3.6. We therefore need to show that if $\pi\neq\mathrm{dec}_{[}n]$, we have $\mathrm{cnat}\left([\right)\pi]<\mathrm{cnat}\left([\right){\mathrm{dec}_{[}n]}]$. For this, note that the permutation graph $G_{\mathrm{dec}_{[}n]}$ associated with the decreasing permutation is the complete graph $K_{n}$ on $n$ vertices, and that $\mathrm{dec}_{[}n]$ is the only permutation whose graph is the complete graph. Now for $\pi\neq\mathrm{dec}_{[}n]$, the permutation graph $G_{\pi}$ is a graph on $n$ vertices which is not complete. The complete graph $K_{n}$ can therefore be obtained from $G_{\pi}$ through a non-empty sequence of edge additions. By Lemma 2.7, this implies that $a_{G_{\pi}}<a_{K_{n}}=a_{G_{\mathrm{dec}_{[}n]}}$. The result then follows from Theorem 2.13. ∎
@@
Theorem 4.23 implies that b(n,(n1)!)=1b(n,(n-1)!)=1. We have also shown (Corollary 4.5) that $b(n,2)=b(n,2!)=(n-2)\cdot 2^{n-3}$, and from [11, Corollary 3.12] we know that $b(n,1)=b(n,1!)=2^{n-2}$. Finally, it was also conjectured in [11] that $b(n,6)=b(n,3!)=\frac{(n-2)(n-3)}{2}\cdot 2^{n-4}$. This suggests that in general, we might have $b(n,k!)=\binom{n-2}{k-1}\cdot 2^{n-1-k}$ for all $n\geq 2$ and $1\leq k<n$ (Sequence A038207 in [16]). However, this formula in fact breaks down almost immediately after these initial values, since our simulations showed that $b(6,24)=31$, which is actually a long way off the value of $\binom{4}{3}\cdot 2^{1}=8$ given by the above sequence. In fact, the sequence $\big(b(n,24)\big)_{n\geq 5}$ does not appear at all in the OEIS, even given just its first three terms $1,31,176$. Table 1 below gives the first few rows of the triangular sequence $\big(b(n,k!)\big)_{n\geq 2,\,1\leq k<n}$.
@@ 5. Conclusion and future perspectives
In this paper, we deepened the combinatorial studies of CNATs initiated in the literature (see e.g.  [3, 11, 2]) by exploiting a connection between CNATs with a given permutation and the acyclic orientation number of the corresponding permutation graph that appeared (in equivalent form) in [13]. In Theorem 3.6, we provided a new bijection between permutations and so-called upper-diagonal CNATs (CNATs whose associated permutation is the decreasing permutation $n(n-1)\cdots 1$), which has the added benefit of preserving certain statistics of these objects. This bijection is defined recursively, based on two operations on labelled CNATs called top-row decomposition and top-row deletion.
We then investigated the enumeration of permutations with a given number of CNATs, answering a number of conjectures from [11]. We gave separate characterisations of permutations with exactly $1$, 22 or $3$ CNATs in terms of so-called quadrants of the permutation’s plot, and in terms of permutation patterns. This allowed us to establish bijections between marked permutations with a single CNAT and permutations with $2$ CNATs (Theorem 4.3), and between permutations with $2$ CNATs and permutations with $3$ CNATs (Theorem 4.15). From this we obtained enumerative formulas for permutations with $2$ or $3$ CNATs (Corollaries 4.5 and 4.16). Finally, we showed that there are no permutations of any length with exactly $5$ CNATs (Theorem 4.17), and that the maximal number of CNATs associated with a given permutation of length $n$ is (n1)!(n-1)!, achieved uniquely for the decreasing permutation $n(n-1)\cdots 1$ (Theorem 4.23).
@@
  • Find a more “direct” combinatorial proof of the fact that there are no permutations with exactly $5$ CNATs, as explained in Problem 4.22.
  • Investigate the enumerative sequences $\big(b(n,k)\big)_{n\geq 1}$ for more values of $k$. In this paper, we have looked at the cases k=1,2,3,5,(n1)!k=1,2,3,5,(n-1)!. The paper [11] suggests that the entries for $k=4,6,7$ do indeed appear in the OEIS [16]. As explained in Remark 4.24, this is not the case for general factorial values of $k$, so it may be that these specific values essentially rely on there being relatively few permutation graphs with the specified acyclic orientation number. For instance, it seems plausible that the only graphs with acyclic orientation number equal to $4$ are the two graphs made up of two triangles in Figures 13 and 18(a) with some trees attached. It also would be particularly interesting to know if there are values of $k$ other than $5$ for which $b(n,k)$ is always 0.
@@
  • A final direction of future research could be to restrict ourselves to certain subsets of permutations, for example derangements (permutations with no fixed point). Fixed points seem to play an important role in the enumeration of permutations according to their number of associated CNATs, since a fixed point implies the existence of a $321$ pattern, which in turn means a $3$-cycle in the permutation graph (and cycles play an important role in graph orientations). For instance, we know that there are no derangements with $2$ CNATs (Proposition 4.9), while permutations with $1$ and $3$ CNATs are all derangements (Propositions 4.2 and 4.14). It is intriguing to consider other values of $k$ and check whether there are derangements or not which have $k$ CNATs. Besides derangements, we could also restrict ourselves to permutations containing or avoiding certain patterns.